I have to solve this problem using the substitution method. 4x-y=6 and 9x-2y=16. I have no idea how to do this.
Answer by ankor@dixie-net.com(22740)

4x - y = 6 and 9x - 2y = 16.
If I do my job right, you will know how to do the substitution method when I am done here
:
First we select which equation to use for substitution, try to choose the one
that has a variable with a coefficient of 1, makes it simpler
:
The 1st equation
4x - y = 6
Subtract 4x from both side
-y = -4x + 6
y has to be positive, multiply each term by -1
y = 4x - 6
:
The 2nd equation:
9x - 2y = 16
we know from the 1st equation that y = (4x-6),
replace y with (4x-6) in the above the above equation
9x - 2(4x-6) = 16
multiply what's inside the brackets by -2 and you have
9x - 8x + 12 = 16
subtract 12 from both sides, combine the x's
x = 16 - 12
x = 4
:
We can use y = (4x-6) to find y, replace x with 4
y = 4(4) - 6
y = 16 - 6
y = 10
:
Our solution: x = 4; y = 10
:
:
Check our solution in the original 2nd equation 9x - 2y = 16, replace x & y
9(4) - 2(10) = 16
36 - 20 = 16; confirms our solutions
:
If you still cannot get your mind around the substitution method, I suggest you
print this out and endeavor to understand each step, individually. it will be
time well spent. C